HETAV PATEL

Grade 11,

5^x+5^(x+1)=125 wat is x value in the given question

5^x+5^(x+1)=125
wat is x value in the given question
 

Grade:12th pass

3 Answers

utpal raj
7 years ago
5^x+5^(x+1)=125
take log on both side 
xlog5+(x+1)log5=log125
(2x+1)log5=log125
2x+1=log125/log5
2x+1=3
so 2x=2
so x=1
should be the ans
hope you understand the procejure.
 
g v subba reddy
7 years ago
taking log  base5 on both sides we will get equation as
log5x +log5x+1 =log125
now the equation can be written as xlog5+xlogx+1=3log5
by cancelling log on both sides we will get x+x+1=3
2x+1=3
2x=2
x=1
Abhishek kumar
6 years ago
Sol:- 5^x+5^(x+1)=125 5^x+x+1=125. (a^m×a^n=a^m+n)5^2x+1=5^3 . (125= 5×5×5=5^3) 2x+1=3. ( By dividing 5 in both term)2x=3-12x=2. (By dividing 2 in both term)So, the value of x = 1